Calculate effective nuclear charge of Fe for 4s electron

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Kajal
 
The atomic number of Fe is 26,
 and thus the nuclear charge of Fe is 26.
The electron configuration of Fe : 1s² 2s² 2p⁶ 3s² 3p⁶ 3d⁶ 4s²
3d orbital is in the 3rd shell, and thus the inner shells including the 1st and 2nd shells (1s² 2s² 2p⁶).
Number of electrons in the inner shells = 2 + 2 + 6 = 10
Effective charge of Fe for 3d electron = 26 - 10 = 16
4s orbital is in the 4th shell, and thus the inner shells including the 1st, 2nd and 3rd shells (1s² 2s² 2p⁶ 3s² 3p⁶ 3d⁶).
Number of electrons in the inner shells = 2 + 2 + 6 + 2 + 6 + 6 = 24
Effective charge of Fe for 4s electron = 26 - 24 = 2